OpenStudy (anonymous):

three-two times V plus 2 =20

OpenStudy (swissgirl):

\({3 \over 4}*V+2=20\)

OpenStudy (anonymous):

20-2=18 18*4=72 72/3=24

OpenStudy (swissgirl):

\({3 \over 4}*V+2-2=20-2\) \({3 \over 4}*V=18\) \({3 \over 4}*{4 \over 3}*V=18*{4 \over 3}\) \(V=24\)